Solve the following system of equations.
x - y = 10
2x + y = 2

Explain which method you chose to solve the system and why you felt it was the best choice

Answer by checkley79(3341)

I'LL CHOOSE THE ELIMINATION METHOD.
LEAST AMOUNT OF WORK. (11 STEPS)
x - y = 10
2x + y = 2 ADD
-------------
3X=12
X=12/3
X=4 ANS.
4-Y=10
-Y=10-4
-Y=6
Y=-6 ANS.
PROOF:
2*4+(-6)=2
8-6=2
2=2
-----------------------
SUBSTITUTION METHOD. (15 STEPS)
X=10+Y
2(10+Y)+Y=2
20+2Y+Y=2
3Y=2-20
3Y=-18
Y=-18/6
Y=-6 ANS.
X-(-6)=10
X+6=10
X=10-6
X=4 ANS.
PROOF:
2*4+(-6)=2
8-6=2
2=2